# Value Gradients with Action Adaptive Search Trees in Continuous (PO)MDPs

### v1

@@ Abstract
Solving Partially Observable Markov Decision Processes (POMDPs) in continuous state, action and observation spaces is key for autonomous planning in many real-world mobility and robotics applications. Current approaches are mostly sample based, and cannot hope to reach near-optimal solutions in reasonable time. We propose two complementary theoretical contributions. First, we formulate a novel Multiple Importance Sampling (MIS) tree for value estimation, that allows to share value information between sibling action branches. The novel MIS tree supports action updates during search time, such as gradient-based updates. Second, we propose a novel methodology to compute value gradients with online sampling based on transition likelihoods. It is applicable to MDPs, and we extend it to POMDPs via particle beliefs with the application of the propagated belief trick. The gradient estimator is computed in practice using the MIS tree with efficient Monte Carlo sampling. These two parts are combined into a new planning algorithm Action Gradient Monte Carlo Tree Search (AGMCTS). We demonstrate in a simulated environment its applicability, advantages over continuous online POMDP solvers that rely solely on sampling, and we discuss further implications.
@@ 1 Introduction
Figure 1: Action adaptive MIS tree with two levels of actions and children state nodes. Usual states for MDPs are replaced by belief states for POMDPs. The actions in the action nodes are updated as the search progresses, indicated with a darker shade, and new states are sampled from the newest actions, denoted in the same color. The MCTS selection path, highlighted in blue, chooses at action nodes the latest action, and then proceeds at random to a child state.
The Markov Decision Process (MDP) and the Partially Observable Markov Decision Process (POMDP) are flexible mathematical frameworks for modeling decision-making under uncertainty. Computing policies for (PO)MDPs that are in continuous in states, actions and observations, has mostly focused on global search algorithms based on POMCP [?], like POMCPOW and PFT-DPW [?], that explore the action space only by sampling. Several algorithms have been suggested for mixed continuous settings, in which not all the state, action and observation spaces are continuous. For continuous state and action spaces, GPS-ABT [?] and LCEOPT [?] use generalized pattern search and cross-entropy optimization respectively, both of which are gradient-free optimization methods. For continuous state and observation spaces, AdaOPS [?] suggested belief packing to combine values for similar beliefs, and LABECOP [?] introduced lazy belief extraction to increase belief diversity.
The first example of an algorithm with performance guarantees for continuous POMDPs was VOWSS [?], that combined Voronoi Optimistic Optimization [?] for action sampling with Weighted Sparse Sampling [?]. The authors also suggested a practical version of the algorithm, VOMCPOW, that combines MCTS with Double Progressive Widening (DPW). [?] presented hierarchical Voronoi-based action sampling, showing improved performance at action dimensionality of 6 to 12.
Our research is motivated by the idea that in many continuous settings, gradient-based optimization is often much faster to converge than gradient-free optimization. It might be prohibitively expensive to sample enough actions to statistically find a near-optimal one. In our approach we consider non-parametric beliefs, making it applicable to a wide range of problems, specifically not restricted by Gaussian or other parametric assumptions. Our methodology consists of two parts. First, we consider how to update actions during search time, without losing information of previous action-value estimates. Second, we derive value gradients w.r.t. actions, for the purpose of computing action gradient updates. We hypothesize that incorporating gradient optimization into continuous (PO)MDP planning will explore the action space more efficiently, especially at high dimensional settings.
A similar idea has been recently explored in the algorithm VG-UCT by [?], where the local gradient-based optimization is combined with the global-search of sampled-based MCTS, but with several drawbacks compared to our approach. Their gradient computation approach is computationally expensive due to being based on finite differences and being recursive. They assume a deterministic transition model given a noise parameter, which is problematic for deriving gradients in POMDPs. Lastly, they do not address the distribution shift that occurs from action updates, which affect the action-value estimates.
Our approach requires access to the state transition model’s likelihood. This is in contrast to previous online (PO)MDP solvers and model-free approaches that consider only use of a generative model. Yet, recent advances in physical simulators have included scalable differentiable physics simulators [??], and GPU-parallelization that have allowed generating accurate simulations with speedups in the scale of orders of magnitudes [??]. We foresee that this will allow to compute the differentiable physical transition models more easily in the near future.
1.1 Additional Related Work

### v3

# Action-Gradient Monte Carlo Tree Search for Non-Parametric Continuous (PO)MDPs

@@ Abstract
Autonomous systems that operate in continuous state, action, and observation spaces require planning and reasoning under uncertainty. Existing online planning methods for such POMDPs are almost exclusively sample-based, yet they forego the power of high-dimensional gradient optimization as combining it into Monte Carlo Tree Search (MCTS) has proved difficult, especially in non-parametric settings. We close this gap with three contributions. First, we derive a novel action-gradient theorem for both MDPs and POMDPs in terms of transition likelihoods, making gradient information accessible during tree search. Second, we introduce the Multiple Importance Sampling (MIS) tree, that re-uses samples for changing action branches, yielding consistent value estimates that enable in-search gradient steps. Third, we derive exact transition probability computation via the Area Formula for smooth generative models common in physical domains, a result of independent interest. These elements combine into Action-Gradient Monte Carlo Tree Search (AGMCTS), the first planner to blend non-parametric particle search with online gradient refinement in POMDPs. Across several challenging continuous MDP and POMDP benchmarks, AGMCTS outperforms widely-used sample-only solvers in solution quality.
@@ 1 Introduction
Planning under uncertainty is a fundamental problem in artificial intelligence, and doing so efficiently holds key for integrating autonomous agents in our daily lives. Specifically in robotics and physical systems, the planning problem is often continuous in states, actions and observations. The Markov Decision Process (MDP) and the Partially Observable Markov Decision Process (POMDP) are flexible mathematical frameworks for modeling long-term decision-making under uncertainty, in which policies are computed to maximize long-term goals. Efficiently solving these problems however is not trivial, with specific research on the continuous case being fairly recent. Online planners mostly focused on exploring the search space by gradually increasing a sampling budget in a technique known as Double Progressive Widening (DPW) [1], with notable POMDP planners being POMCPOW and PFT-DPW [2]. This blind approach is often inefficient, as it requires evaluating many actions to find a near-optimal one. To address this, recent work in continuous domains has explored black-box optimization methods that guide sampling. Techniques such as Voronoi optimistic optimization [3, 4, 5], kernel regression [6] and Bayesian optimization [7, 8] have all been proposed to bias action selection towards promising regions. However, they do not leverage the internal structure of the transition or observation model, missing the opportunity for model-aware introspection to improve planning efficiency.
Our research is motivated by the idea that in many continuous settings, gradient-informed optimization is often much faster to converge than gradient-free methods. It might be prohibitively expensive to sample enough actions to statistically find a near-optimal one. Yet, computing the gradient of the action-value function is not trivial. Recently Lee et al. [9] explored gradient steps integrated in MCTS for continuous MDPs. Their analysis was limited to Lipschitz-continuous transition models and rewards, they did not analyze the value estimate drift that occurs from action updates, and most importantly their derivation does not extend to POMDPs trivially. Our gradient results can be seen as similar to the policy gradient theorem [10], however we calculate score gradients w.r.t. actions, rather than policy parameters, and in POMDPs. Policy gradients in POMDPs have been limited to memory-less policies [11] or to finite action spaces [12], yet we assume belief-dependent policies.